Identify the following components of each quadratic function.
Then graph the function.
1. y=x^2+2x-3
axis of symmetry x=-b/2a :
Vertex (-b/2a,y):
y-intercept (c):
two other points:
help please, show all work thanks so much!

Answer by ewatrrr(24785)

 
Hi
y = x^2 + 2x - 3 = (x+1)^2 - 1 - 3 = (x+2)^2 - 4
axis of symmetry x=-b/2a : x = -1
Vertex (-b/2a,y): (-1,-4)
y-intercept (c): y = -3
two other points: (1,0) and (-1, -4)